OpenStudy (anonymous):

solve what are the solutions and find the x-intercepts : 7x+x(x-4)=0 solve what are the solutions and find the x-intercepts : 7x+x(x-4)=0 @Mathematics

OpenStudy (across):

\[x^2-4x+7x=0\]\[x^2+3x=0\]\[x^2+3x+\frac{9}{4}=\frac{9}{4}\]\[\left(x+\frac{3}{2}\right)^2=\frac{9}{4}\]\[x+\frac{3}{2}=\pm\frac{3}{2}\]\[x=0,3\]

OpenStudy (anonymous):

x = 0 and x = 3

OpenStudy (anonymous):

Ah across gotcha.

OpenStudy (anonymous):

The solution set is {-3, 0} (x=3 is not a solution.)

OpenStudy (across):

Mandolino is right. Good observation.

OpenStudy (asnaseer):

@across, when you got to:\[x^2+3x=0\]this could have been simplified to:\[x(x+3)=0\]so x=0 or -3

OpenStudy (anonymous):

Another method: you could factor out an x 7x+x(x-4)=0 x(7+x-4)=0 x(x+3)=0 x=0 or x+3=0 x=-3 Just for the sake of interest :)

OpenStudy (anonymous):

yeah I mess that up, x does not equal 3.

OpenStudy (anonymous):

so the solutions are 0 and 3 and the x intercepts would be 3,0 and 0,0???

OpenStudy (anonymous):

No, the solutions are 0 and -3

OpenStudy (anonymous):

o thanks i see it now

OpenStudy (anonymous):

The solutions to the equation set equal to 0 are also the x-intercepts if we think of it as a function.

OpenStudy (anonymous):

does anyone know how to find and label the vertex and line of symmetry?